# MIRI spectrophotometry of GN-z11: Detection and nature of an optical red continuum component

### v1

@@ Acknowledgements.
As mentioned in Sect. 3.1, we used the NIRCam photometry to renormalise the NIRSpec spectra. In this appendix we introduce the main steps followed during this renormalization.
@@
First, we PSF-matched the NIRCam images to the MIRI/F560W resolution. This is done by deriving empirical PSFs (ePSFs) based on the stars available in the FoV (i.e., between 3 and 6 non-saturated in each filter). We then create the corresponding kernel for each NIRCam image using the MIRI PSF models from Libralato+24, as there are not enough available stars in the MIRI FoV. Once the convolution has been applied, we extracted the NIRCam and MIRI photometry using circular apertures of = 0.5″(see Sect. 1 for details). As a consistency check, we extracted the photometry at the native spatial resolution of each filter, applying the corresponding aperture corrections. The resulting fluxes are consistent with those obtained from the PSF-matched data within uncertainties.
@@
We noticed that our NIRCam photometry is slightly larger (20 25) than the one derived in Tacchella+23_GNz11 for a smaller aperture ( = 0.35″). In our analysis, we have used a more recent data release from JADES (i.e., DR3) where the NIRCam images were reduced with improved calibration files from JWST. In addition, we have used a larger and more distant from GN-z11 annular aperture to derive the background level. The combined effect of using a more distant aperture to derive the background level and more recent calibration files and pipeline favour the detection of diffuse emission that was missed in the original calibrated images.
@@
Finally, we convolved the NIRSpec/Prism spectra with the NIRCam filter transmission curves to derive the expected fluxes in each filter. These NIRSpec continuum values were therefore compared with the NIRCam photometry, measuring a median 32.5 offset that remains nearly constant with wavelength, showing variations of less than 5. We had also compared the NIRCam photometry with the NIRSpec pseudo-continuum fluxes obtained in 750 bins (see Sect. 3.3), obtaining similar results. We therefore multiply the NIRSpec/Prism spectra by 1.325 to match the NIRCam and NIRSpec data. Bunker+23 reported an agreement between the NIRSpec/Prism reduction, which included slit-loss corrections for point-like source, and the = 0.1″ NIRCam photometry from Tacchella+23_GNz11. This difference is produced by as GN-z11 UV morphology deviates from an ideal point-like source, which is assumed during the data reduction. In fact, even larger differences between NIRCam and NIRSpec/MSA (i.e., 30 50) have been found in other high- galaxies (e.g., JADES-GS-z14-0, Carniani+24).
@@
For consistency, we tested whether this difference in the UV flux would lead to any noticeable difference during the SED fitting. Therefore, we replicated the SED analysis described in Sect. 4.3, with a double stellar population and the presence of an AGN, considering the NIRSpec pseudo-filters prior normalization. As in this work we assumed a uniform re-normalization, most of the properties derived with the UV slope remains unaltered. However, having a lower UV continuum results in a slightly older (i.e., 6 Myr vs 4 Myr) and more massive young stellar population (i.e., 13 larger). As the H and the optical and near-IR fluxes do not change, we do not observe significant differences either in the SFR, the total stellar mass or the AGN luminosity.

### v2

@@ Appendix A NIRCam-NIRSpec-MIRI normalisation
As mentioned in Sect. 3.1, we used the NIRCam photometry to re-normalise the NIRSpec spectra. In this appendix we introduce the main steps followed during this re-normalisation.
@@
First, we PSF-matched the NIRCam images to the MIRI/F560W resolution. This is done by deriving empirical PSFs (ePSFs) based on the stars available in the FoV (i.e. between 3 and 6 non-saturated in each filter). We then create the corresponding kernel for each NIRCam image using the MIRI PSF models from Libralato et al. (2024), as there are not enough available stars in the MIRI FoV. Once the convolution has been applied, we extracted the NIRCam and MIRI photometry using circular apertures of = 0.5″(see Sect. 1 for details). As a consistency check, we extracted the photometry at the native spatial resolution of each filter, applying the corresponding aperture corrections. The resulting fluxes are consistent with those obtained from the PSF-matched data within uncertainties.
@@
We noticed that our NIRCam photometry is slightly larger (20 25) than the one derived in Tacchella et al. 2023a for a smaller aperture ( = 0.35″). In our analysis, we have used a more recent data release from JADES (i.e. DR3) where the NIRCam images were reduced with improved calibration files from JWST. In addition, we have used a larger and more distant from GN-z11 annular aperture to derive the background level. The combined effect of using a more distant aperture to derive the background level and more recent calibration files and pipeline favour the detection of diffuse emission that was missed in the original calibrated images.
@@
Finally, we convolved the NIRSpec/Prism spectra with the NIRCam filter transmission curves to derive the expected fluxes in each filter. These NIRSpec continuum values were therefore compared with the NIRCam photometry, measuring a median 32.5 offset that remains nearly constant with wavelength, showing variations of less than 5. We had also compared the NIRCam photometry with the NIRSpec pseudo-continuum fluxes obtained in 750 bins (see Sect. 3.3), obtaining similar results. We therefore multiply the NIRSpec/Prism spectra by 1.325 to match the NIRCam and NIRSpec data. Bunker et al. (2023) reported an agreement between the NIRSpec/Prism reduction, which included slit-loss corrections for point-like source, and the = 0.1″ NIRCam photometry from Tacchella et al. 2023a. This difference is produced by as GN-z11 UV morphology deviates from an ideal point-like source, which is assumed during the data reduction. In fact, even larger differences between NIRCam and NIRSpec/MSA (i.e. 30 50) have been found in other high- galaxies (e.g. JADES-GS-z14-0, Carniani et al. 2024).
@@
For consistency, we tested whether this difference in the UV flux would lead to any noticeable difference during the SED fitting. Therefore, we replicated the SED analysis described in Sect. 4.3, with a double stellar population and the presence of an AGN, considering the NIRSpec pseudo-filters prior normalisation. As in this work we assumed a uniform re-normalisation, most of the properties derived with the UV slope remains unaltered. However, having a lower UV continuum results in a slightly older (i.e. 6 Myr vs 4 Myr) and more massive young stellar population (i.e. 13 larger). As the H and the optical and near-IR fluxes do not change, we do not observe significant differences either in the SFR, the total stellar mass or the AGN luminosity.